OpenStudy (anonymous):

Forty-two is what percent of 70? A. 21% B. 60% C. 74% D. 167%

OpenStudy (anonymous):

B

OpenStudy (anonymous):

are you right

OpenStudy (anonymous):

70*0.60=42

OpenStudy (k_lynn):

YOu need to set up a proportion. Since percents are out of 100, the proportion should look like this. \[\frac{ 42 }{ 100 }=\frac{ ? }{ 70 }\] Now all you need to do is cross multiply 42 and 70. then divide by 100.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get 29.4

OpenStudy (k_lynn):

hmm you're right. And that isn't an option. @iGreen can you tell me if I screwed up or not?

OpenStudy (anonymous):

i got 60

OpenStudy (anonymous):

@iGreen

OpenStudy (aaronandyson):

take it as 42 = x/70 times 100 yet its same ?

OpenStudy (aaronandyson):

42/70 = 6/10 = 60/100 = 60%

OpenStudy (anonymous):

im confussed

OpenStudy (bibby):

you can just set it up as an equation and find the decimal value, then convert to percent Forty-two is what percent of 70? 42=70x x=42/70

OpenStudy (anonymous):

@k_lynn had the right idea but wrong ratios $$\frac{x}{100} = \frac{42}{70}$$

OpenStudy (anonymous):

easiest way to do it imo, is to take 42/70 = 0.6, then multiply by 100

OpenStudy (anonymous):

as @bibby mentioned

OpenStudy (anonymous):

i got 60

OpenStudy (anonymous):

looks good to me